Evaluate each function for the value specified. f(x) = x^2 - x -2; find f(-2)
Answer by checkley77(12844)

f(x) = x^2 - x -2; find f(-2)
f(-2)=-2^2-(-2)-2
f(-2)=4+2-2
f(-2)=4 ans.